What is the total charge of 100kg of electrons?


1
Expert's answer
2021-04-20T16:50:02-0400

The following amount of electrons is contained in the mass of "m = 100kg":


"N = \dfrac{m}{m_e}"

where "m_e = 9.11\times10^{-31}kg" is the mass of one electron. The charge that is carried by this number of electrons is given as follows:


"q = Ne = \dfrac{me}{m_e}"

where "e = -1.6\times 10^{-19}C" is the charge of one electron. Thus, obtain:


"q = \dfrac{100\cdot (-1.6\times 10^{-19}) }{ 9.11\times10^{-31}} \approx -1.76\times10^{13}C"

Answer. "-1.76\times10^{13}C".